# A Review of Community-Centric Power Systems Resilience Assessment and Enhancement Strategies

### v1

@@ 6 Techno-Legal Aspects of Resilient Power Systems and Communities
The enhancement of power systems resilience, as reviewed in this paper, transcends overwhelmingly technical implementation. This focus is critical, as power systems function as the keystone infrastructure within a complex SoS. This dependency elevates the power grid to a primary technical and regulatory priority, as a failure cascade originating from the grid can paralyze an entire community. This section, therefore, bridges the resilience assessment methods, AI applications, and interdependency models discussed previously with the evolving transatlantic legal regimes. We analyze the implications for system operators, municipalities, and community energy groups, and align these legal considerations with the paper’s focus. Figure 7 conceptualizes the power grid as the central keystone within a SoS architecture. The diagram illustrates how the techno-legal frameworks (e.g.North American Electric Reliability Corporation Critical Infrastructure Protection -NERC CIP- [104], NIS2 Directive [105]) impose a top-down regulatory structure (Tier 1) on the power system (Tier 2). This governance is a direct consequence of the bottom-up dependency (Tier 3), where all other critical societal infrastructures (healthcare, water, transport) fundamentally rely on the power system for their own operation.
@@
6.1 The Evolving EU Regulatory Landscape
The relevant European Union (EU) regulatory landscape is multifaceted. The GDPR establishes a strict regime for processing personal data, which is pertinent to outage analytics, Advanced Metering Infrastructure (AMI), and the use of social vulnerability inputs [106]. Concurrently, the NIS2 Directive imposes comprehensive risk management, supply chain security, and incident reporting obligations on essential entities within the energy sector [105]. This is complemented by the CER Directive, which mandates risk assessments and resilience measures to protect critical infrastructure. Finally, the forthcoming EU AI Act introduces a risk-based framework for AI systems, classifying safety components in critical infrastructure and AI models that affect grid services as high-risk, thereby imposing stringent duties regarding data quality, risk management, transparency, and human oversight.
6.2 Comparative US Regulatory Context
In the United States, a parallel framework exists, though it is often more sectoral. Mandatory cybersecurity standards for the bulk electric system are set by the NERC Critical Infrastructure Protection (CIP) standards, enforced by Federal Electric Regulatory Console (FERC) [104]. Broader critical infrastructure guidance and risk management frameworks are promulgated by the Cybersecurity and Infrastructure Security Agency (CISA) and the National Institute of Standards and Technology (NIST), particularly the widely adopted NIST Cybersecurity Framework (CSF)[107]. For data privacy, the US lacks a single federal equivalent to GDPR, relying instead on a patchwork of state-level laws (e.g. the California Consumer Privacy Act - CCPA) [108] and federal sectoral laws (e.g. Health Insurance Portability and Accountability Act of 1996, which becomes relevant when correlating outages with vulnerable health [109]). For AI governance, the US is currently focused on non-binding frameworks, notably the NIST AI Risk Management Framework (AI RMF) [110], which encourages a voluntary, evidence-based approach to managing AI risks, contrasting with the EU’s binding, compliance-driven AI Act [112].
6.3 Implications for Data and Models in Resilience Analytics
These legal frameworks have direct implications for the data and models utilized in resilience analytics. The processing of social and outage data, including health, demographic, or social vulnerability features, demands strict purpose limitation, aggregation, or the application of differential privacy under GDPR[106], or adherence to de-identification standards under HIPAA [109]. Privacy-preserving learning methodologies are preferable when correlating outage data with vulnerable populations, directly supporting the community-centric metrics reviewed. Furthermore, asset and interdependency data, such as grid topology, substation layouts, and linkages to gas, water, or transport systems, must be treated as sensitive operational data under NERC CIP [104] and NIS2[105]. Where real topologies are restricted for security reasons, validated synthetic networks should be employed to support the fragility-based and data-driven assessments.
6.4 Cybersecurity by Design for Resilient Operation
Achieving resilience necessitates a cybersecurity-by-design approach, integrating controls from frameworks like NIS2 [105] and NERC CIP [104] directly into resilience functions. Technical workflows for defensive islanding, DER dispatch, and network reconfiguration must be mapped to specific controls such as robust authentication, least privilege access, network segmentation, and secure update protocols. This extends to the secure supply chain for operational technology, including IEDs and PMUs, where requirements for Software Bills of Materials (SBOMs) [111] and vulnerability disclosure are essential. Incident response procedures must also be designed for cyber-physical scenarios, particularly those that degrade communications during critical operations like MG formation.
6.5 Law-Tech Integration in Interdependent Infrastructures
The techno-legal framework must directly address the SoS realities of interdependent infrastructures. Resilience planning should include joint exercises for power, gas, water, and health services, supported by data sharing templates that pre-define lawful bases (under GDPR) and confidentiality clauses, reflecting the cascading risks detailed in this review. In procurement and liability, tenders for grid equipment must embed relevant clauses (e.g., NIS2 [105], NERC CIP [104]conformity, AI Act-ready documentation [112]), requiring evidence of secure-by-design principles. For AI-assisted decisions, liability allocation must be supported by immutable logs linking sensor inputs, model outputs, and human approvals.
To ensure accountability, techno-legal performance indicators—such as time to detect and report cyber incidents or percentage of datasets with completed Data Protection Impact Assessments (DPIAs), should be tracked with the same rigor as traditional resilience indices [127].
7 Research Gaps, and Future Research Directions
Despite the development of various metrics to evaluate power system resilience and subsequent enhancement strategies, current methodologies have the following limitations.
  1. Comprehensive Failure Models: Current models like fragility curves do not adequately capture the spatiotemporal dynamics of natural disasters or their impact on grid components. Improved fragility curves, alongside scenario-based simulations, could lead to more accurate failure models [11].
  2. Data Utilization and AI/ML Algorithms: The advancement of AI has significantly enhanced the ability to process large datasets for forecasting outages and assessing resilience. However, the scarcity of historical data and the challenges of data collection remain significant barriers. There is a crucial need for collaboration between utilities and researchers to overcome these obstacles and for effective data integration and cleansing practices to utilize big data fully [29].
  3. Confidentiality and Data Sharing: The confidentiality surrounding vital infrastructure data often limits the availability for research. The real topology of power networks is often proprietary and highly sensitive. However, synthetic networks can be generated using publicly available datasets, which share similar properties with real network topologies [36].
  4. Extreme Weather Events Modeling: Due to the rarity of extreme weather events, it is difficult to collect adequately large extreme weather event datasets. With generative AI, extreme weather scenario data can be generated, which shares a similar distribution to the real extreme weather events [37]. Thus, the resilience of the power systems can be simulated before such events even happen in the real world. Planning for the recurrence of extreme events shortly after an initial event is identified as a key priority for future research.
  5. Integration of AL/ML with Stochastic Methods: Combining AL and ML techniques paired with stochastic approaches could significantly improve the modeling of nonlinear behaviors in power systems. This integration promises more precise predictions and assessments, particularly beneficial for complex and large-scale systems.
8 Conclusion
This paper offers a comprehensive review of community-centric power system resilience assessment, summarizing frameworks, characteristics, and enhancements. Resilience assessment and enhancement play a vital role in facilitating resilient planning and operation. The paper identifies two fundamental approaches to evaluating power system resilience. The first is detailed engineering: this approach focuses on developing the fragility curves of the components within a system to assess resilience. The second is data-driven: this method relies on aggregated outage data to evaluate resilience at a broader level, using frameworks such as the resilience triangle, resilience trapezoid, or resilience curves. The community-centric resilience assessment necessitates considering socioeconomic and demographic factors, as well as the interdependence between power and other energy systems. Next, the paper presents a comprehensive review of numerous research papers concerning the improvement of resilience within power systems. Moreover, the application of AI and ML technologies in power system resilience is explored, which facilitates the advanced analysis of system responses to extreme events.
Beyond technical metrics, this review shows that power systems sit at the center of a techno legal system of systems, where frameworks such as NERC CIP, NIS2, GDPR, and the EU AI Act shape how data, AI tools, and digital twins support resilience. Embedding these obligations into grid design, operation, and AI-driven decision support strengthens community-centric resilience while aligning outage analytics, interdependency modeling, and cyber-physical protection with legal accountability. Future work should integrate resilience indices with compliance-oriented indicators, linking system performance to measurable duties for operators and for other critical infrastructures that depend on the grid.

### v2

@@ 5 Techno-Legal Aspects of Resilient Power Systems and Communities
Community-centric power system resilience is not determined by technical design alone. It also depends on how data are collected, protected, and governed across utilities, healthcare providers, and other critical infrastructures. This becomes particularly important as resilience enhancement strategies increasingly rely on AI-enabled outage prediction, automated restoration support, DER coordination, network reconfiguration, and the operation of resilience hubs. In these settings, legal and regulatory frameworks do not merely constrain implementation after engineering decisions have been made; they directly influence what data may be used, which models are deployable, and how accountability is assigned when failures affect vulnerable communities.
Accordingly, techno-legal governance should be treated as an integral component of community-centric resilience, rather than a downstream compliance layer. As such, resilience evaluation must also account for privacy, cybersecurity, equity, traceability, and lawful coordination across infrastructures and communities. Figure 6 provides a system-level representation of these interactions, illustrating how community needs and critical services define resilience priorities and drive the collection of diverse data inputs. These inputs include advanced metering infrastructure (AMI) and smart meter data, outage histories, weather and hazard data, topology and asset condition information, and social vulnerability indicators. Such data streams enable AI-driven resilience functions.
Within this architecture, techno-legal governance mechanisms, such as NERC CIP standards [76], the Network and Information Security Directive 2 (NIS2) [37], and data protection regulations including GDPR, the Health Insurance Portability and Accountability Act (HIPAA), and the California Consumer Privacy Act (CCPA)—act as cross-cutting constraints across all layers. These frameworks govern data collection and sharing, regulate AI model design, enforce cybersecurity and operational compliance, and establish accountability, auditing, and liability mechanisms. These governance requirements arise from the strong interdependencies among critical infrastructures. As a result, resilience must be understood as a coordinated, community-centric objective shaped by infrastructure interdependencies, data-driven decision-making, and regulatory constraints.
@@
5.1 Data Governance and Privacy Constraints in Resilience Analytics
The AI- and data-driven methods rely on the integration of diverse datasets, including outage histories, grid topology, and weather data. At the community level, resilience-oriented analytics may also require demographic, socioeconomic, and health-related information to identify populations that are likely to experience disproportionate harm during prolonged outages. This creates a central tension for community-centric resilience planning. The models that appear most accurate often depend on fine-grained and heterogeneous data, yet the collection and fusion of such data are restricted by privacy and data governance regimes.
Under the GDPR, the processing of personal data must satisfy principles such as lawfulness, purpose limitation, data minimization [36]. These principles are directly relevant when outage prediction or restoration prioritization models incorporate household-level consumption records, location-linked vulnerability indicators, or data associated with medically dependent users. In the United States, similar concerns arise through sectoral and state-specific frameworks, including HIPAA for health-related information [105] and consumer privacy laws such as the California Consumer Privacy Act [22]. For community-centric resilience, the practical issue is not only whether data are technically available, but whether their integration is legally justified.
This tension is especially important for models that aim to identify vulnerable communities or prioritize resilience investments. As discussed in Section 4, outage impacts vary significantly across communities depending on social vulnerability. However, identifying a “vulnerable” customer or neighborhood requires the processing of granular data related to health, income, disability, age, or other socially sensitive attributes. This means that resilience planning increasingly falls within the scope of strict privacy and data governance obligations. In practice, the stronger the ambition to make resilience more targeted and equitable, the greater the need for governance mechanisms that lawfully manage sensitive data.
The same issue extends to the AI methods reviewed earlier. High predictive performance alone is not sufficient to justify the operational use of a model in resilience planning. For example, outage prediction systems trained on highly granular household and infrastructure data may improve forecasting performance, but they may also become difficult to reconcile with data minimization requirements, purpose limitation, or restrictions on cross-sector data sharing. In addition, black-box models with large data requirements may reduce transparency regarding data provenance. This concern is particularly salient for advanced transformer-based and LLM-based approaches discussed in Section 3. Although these models may show strong performance in selected prediction tasks, their large computational requirements, opacity, and dependence on extensive training data make them difficult to explain and validate for critical infrastructure use. Their tendency to produce unreliable or hallucinated outputs further limits their suitability for safety-critical and accountability-sensitive settings.
These constraints imply that future resilience analytics should be designed with privacy preservation as a first-order requirement. Techniques such as data aggregation, differential privacy, and federated or distributed learning are promising because they reduce the need to centralize all sensitive data while still enabling collaborative modeling. Similarly, asset and interdependency data, including grid topology, and substation layouts should be treated as sensitive operational information under frameworks such as NERC CIP [76] and NIS2 [37]. In these cases, validated synthetic test systems and controlled data-sharing mechanisms may be preferable for resilience analysis.
5.2 Cybersecurity by Design for AI-Enabled Resilient Operation
Cybersecurity is not external to resilience engineering. It is a core condition for deploying AI-enabled control and decision-support systems. Several resilience enhancement strategies now depend on digital control, advanced sensing, and learning-based optimization. Examples include repair crew scheduling, MG formation, and network reconfiguration. When these functions are implemented through AI-enabled or data-driven architectures, they inherit not only the benefits of fast adaptation and improved decision support, but also new vulnerabilities related to cyber intrusion, telemetry corruption, and software supply-chain compromise.
This concern is particularly important for RL and deep reinforcement learning (DRL) methods. These methods have been proposed for post-disaster network reconfiguration [130, 49, 2, 48], restoration support [66, 50, 117], and voltage regulation [28]. However, their performance is usually demonstrated under simulated assumptions. In practice, community-centric resilience operations occur under precisely the opposite conditions. Extreme events may degrade communications, distort sensor visibility, interrupt centralized coordination, and create out-of-distribution physical states. Under these conditions, an AI agent may receive manipulated, delayed, or incomplete inputs and may issue unsafe actions.
This creates a direct bridge between resilience enhancement and cybersecurity regulation. NERC CIP [76] and NIS2 [37] both imply that resilience functions must be developed and operated with secure architectures. For AI-enabled resilience functions, they should be assessed for secure sensing, human override, and decision traceability. This is especially important when resilience operations involve PMUs, SCADA, or field devices whose compromise may produce cascading physical effects.
If RL is used for network reconfiguration, then the integrity of topology data and switching status becomes a core safety issue. If AI is used for repair crew scheduling, the model must remain reliable when communication systems are degraded. In this sense, the cyber-resilience of AI-driven functions becomes a research problem in its own right. Future work should therefore move beyond proof-of-concept control performance and address formal verification, adversarial robustness, runtime monitoring, secure-by-design requirements [103], and human-in-the-loop supervisory control for AI-assisted grid operations.
5.3 Law-Tech Integration in Interdependent Infrastructures and Resilience Hubs
As discussed in Section 4, outages propagate through interdependent infrastructures and create disproportionately severe consequences for socially vulnerable populations. This implies that resilience is not only about asset recovery, but also about ensuring lawful, secure, and coordinated action across institutions and sectors.
This point is especially important for resilience hubs. To function effectively, resilience hubs must be integrated into the wider cyber-physical system of grid operations, restoration prioritization, and interdependent service provision. Restoration and scheduling models can prioritize feeders connected to hubs and medically critical facilities. Network reconfiguration and MG formation strategies can be used to preserve or restore electricity supply to hubs during wider service disruption.
This interpretation also raises a set of governance questions. Which entities are lawfully permitted to share data on medically dependent populations or socially vulnerable households. Under which legal basis can certain facilities or neighborhoods be prioritized during restoration. How should decision authority, liability, and accountability be allocated when AI-assisted tools are used. These questions show that community-centric resilience depends on institutional preparedness as much as technical preparedness.
For this reason, law-tech integration in interdependent infrastructures should include predefined data-sharing templates, lawful bases for emergency processing under GDPR [36], and cross-sector emergency exercises that reflect cascading infrastructure risks. Procurement processes for grid and community resilience equipment should embed explicit requirements for NIS2 [37], NERC CIP [76], and AI-governance readiness [39]. In parallel, techno-legal performance indicators, such as time to detect and report cyber incidents, and completeness of cross-sector data governance documentation, should be tracked alongside traditional resilience metrics [53].
5.4 Comparative EU and US Regulations and Their Engineering Implications
The EU and US regulatory landscapes both address critical infrastructure resilience, cybersecurity, and AI oversight, but they do so through different legal structures. These differences have direct consequences for the design, procurement, and operational deployment of community-centric resilience solutions.
In the EU, the regulatory environment is becoming increasingly integrated through the combined effects of the GDPR [36], the NIS2 Directive [37], the CER Directive, the Energy Efficiency Directive [38], and the EU AI Act [39]. Together, these instruments create a structured governance environment in which data protection, cybersecurity, and AI oversight are increasingly linked. This has several technical implications. First, resilience-oriented AI systems that affect critical infrastructure operations are likely to face stronger requirements. Second, the stronger emphasis on accountability may favor more interpretable and auditable AI architectures. Third, compliance requirements are likely to shape model documentation, validation workflows, and post-deployment monitoring.
In the United States, the governance picture is more fragmented. Cybersecurity obligations for the bulk electric system are more mature in some respects through NERC CIP, under the oversight of the Federal Energy Regulatory Commission (FERC) [76]. Broader risk management guidance is also available through CISA and the NIST Cybersecurity Framework [75]. However, there is no single federal equivalent to the GDPR for general personal data protection. Instead, privacy obligations arise through a patchwork of state-level and sector-specific rules. AI governance remains more reliant on voluntary or soft-law frameworks such as the NIST AI Risk Management Framework [74]. This fragmented structure has practical implications for engineering deployment. A resilience-enhancing AI platform that uses customer-level, health-related, or community vulnerability data may be deployable under one set of conditions in one state and under different or less clear conditions elsewhere. This complicates data integration, benchmarking, and procurement across service territories.
For engineering audiences, the key point is that regulatory context influences technical choices. In the EU, stronger and more unified obligations may increase the burden of compliance, but they also provide a clearer direction toward explainable, traceable, and human-supervised resilience systems. In the US, the greater variation across jurisdictions may allow local experimentation, but it may also hinder standardization and equitable scaling of community-centric resilience tools. Therefore, future resilience research should examine how regulatory structures shape data architecture, human oversight mechanisms, and the long-term deployability of AI-enabled resilience strategies.
6 Conclusions and Discussions
This paper presents a comprehensive review of community-centric power system resilience, with a focus on AI-enabled methods, critical infrastructure interdependencies, and techno-regulatory considerations. The central theme of the paper is disconnect between engineering metrics, social vulnerability indices, and regulatory definitions, leading to a fragmented understanding of resilience as a concept. In addition, there is a tension between data-driven personalization and data privacy, which runs throughout both the AI and legal perspectives on resilience assessment and enhancement methods. The paper also introduces resilience hubs as a key example of a comprehensive, community-centered approach to addressing interdependent and multifaceted vulnerabilities.
Existing resilience metrics are predominantly system-centric (e.g., load served or infrastructure performance) rather than community-centric (e.g., impacts on critical social services and vulnerable populations). Many studies continue to evaluate power system resilience independently of interdependent infrastructures, limiting the ability to quantify cascading impacts. Furthermore, most data-driven approaches focus primarily on outage prediction, while comparatively little attention has been given to modeling community recovery trajectories, service restoration prioritization, and adaptive resilience strategies. Social vulnerability dynamics during outages are also often neglected in resilience assessments. In addition, current regulatory frameworks largely emphasize reliability compliance rather than measurable community resilience performance, and community-level resilience studies rarely incorporate cyber disruptions affecting grid operations and DER coordination.
This paper highlights several pressing research gaps in this area. For example: 1) Privacy-preserving federated learning architectures for outage prediction that comply with both GDPR and NERC CIP requirements remain underdeveloped. 2) Formal verification frameworks for ensuring the safety and robustness of DRL-based network reconfiguration agents under cyber-physical attack scenarios, as outlined in NIS2, are still in their early stages of development. 3) Privacy-preserving resilience analytics remain insufficiently developed, especially for outage prediction and restoration prioritization models. 4) AI-enabled operational methods, especially RL-based approaches for DER coordination, and restoration support, require much stronger safety assurance under cyber-physical attack, telemetry disruption, and extreme out-of-distribution conditions. 5) Resilience hubs and other community-critical facilities need governance-ready design frameworks that integrate lawful data sharing, operational prioritization, and cross-sector accountability. 6) The evaluation of resilience-oriented AI models should move beyond predictive accuracy to include explainability, privacy robustness, and regulatory deployability. 7) Comparative regulatory analysis should be used to inform standardization, so that resilience tools are designed for real-world implementation rather than only for benchmark performance.
Taken together, these issues show that community-centric resilience must be framed as a cyber-physical-social governance challenge. The future research agenda should therefore focus on integrated methods that connect engineering performance with social vulnerability, secure AI system design, and legally compliant operational workflows. Such an agenda would better align resilience investments with the needs of vulnerable communities while improving the trustworthiness and deployability of AI-enabled power system resilience solutions.
Acknowledgment
This work is partially supported by the DOE under Award Number DE-EE0010413. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of DOE. The first author acknowledges earlier discussions with Dr. Caisheng Wang on this topic.